# Large Language Models as Natural Selector for Embodied Soft Robot Design

### v1

@@ IV-A Evaluation Metrics
Task Name # Questions Reward Diff Max Reward Min Reward
Pusher-v1 3000 0.04 to 1.39 0.21 -1.33
Jumper-v0 0.13 to 8.07 0.04 -8.44
Carrier-v1 0.90 to 6.62 3.60 -5.85
Walker-v0 0.74 to 2.89 4.09 -1.21
UpStepper-v0 0.08 to 2.51 1.64 -1.27
Balancer-v0 0.12 to 1.05 0.13 -1.41
Climber-v0 0.01 to 0.19 0.16 -0.18
Pusher-v0 0.49 to 3.07 3.07 -1.66
Carrier-v0 0.53 to 3.50 2.81 -2.86
GapJumper-v0 0.08 to 2.62 2.28 -0.77
BridgeWalker-v0 0.39 to 3.69 4.04 -1.49
BidirectionalWalker-v0 0.31 to 4.16 1.81 -3.69
Summary 36000 0.01 to 8.07 4.09 -8.44
Difficulty-Weighted Accuracy (DWA): Quantifies the model’s ability to discern subtle differences in robot task performance, with higher weights assigned to harder questions (smaller reward differences). Let $N_{tasks}$ be the total number of robot tasks. For each robot task $t\in\{1,...,N_{tasks}\}$, define:
$\text{Difficulty}_{t,i}=1-\frac{\Delta P_{t,i}-\Delta P_{t}^{\text{min}}}{\Delta P_{t}^{\text{max}}-\Delta P_{t}^{\text{min}}}$ (3)
$\text{DWA}_{t}=\frac{\sum_{i=1}^{N_{t}}\mathbb{I}(\text{Correct}_{t,i})\times\text{Difficulty}_{t,i}}{\sum_{i=1}^{N_{t}}\text{Difficulty}_{t,i}}$ (4)
Cross-Task Normalization: Aggregates task-specific DWAs while eliminating scale variance:
Global $\text{Global DWA}=\frac{1}{N_{tasks}}\sum_{t=1}^{N_{tasks}}\text{DWA}_{t}$ (5)
where $\Delta P_{t,i}=P_{t,i}^{\text{correct}}-P_{t,i}^{\text{incorrect}}$ is the reward gap between correct and incorrect options for question $i$ in robot task $t$, $\Delta P_{t}^{\text{min}}$ and $\Delta P_{t}^{\text{max}}$ are the minimum and maximum reward gaps within task $t$, $\mathbb{I}(\cdot)$ is the indicator function, and $N_{t}$ is the number of questions per task. Higher DWA indicates better performance on discriminative tasks with small reward differences.
@@ IV-B Benchmarking Results of Robot Design Selection
We conduct a comparative evaluation of four most advanced available Large Language Models (LLMs) across 12 diverse environments, using metrics categorized into ‘Easy’ and ‘Hard’ tasks. The summarized results of accuracy and consistency are presented in Table II.
Category Task Name GPT-o3-mini Gemini-2.0-flash Gemini-1.5-Pro Grok-2
Accuracy Consistency Accuracy Consistency Accuracy Consistency Accuracy Consistency
Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard Easy Hard
Locomotion Walker-v0 57.65 56.58 73.11 72.73 71.40 63.00 94.15 94.75 75.40 65.20 96.10 96.75 71.33 61.70 90.60 89.85
BridgeWalker-v0 49.65 44.86 72.17 77.99 75.40 61.80 95.70 91.55 81.57 68.20 97.95 97.50 78.40 63.67 91.20 89.25
BidirectionalWalker-v0 51.46 52.90 72.91 73.26 55.23 56.11 93.60 95.31 53.97 55.03 97.15 97.00 55.30 53.50 88.20 90.70
UpStepper-v0 63.57 53.72 79.02 78.45 72.97 56.70 95.95 93.95 72.97 61.10 97.00 95.40 74.00 58.43 94.85 92.00
GapJumper-v0 58.88 53.63 77.62 76.31 71.53 58.83 94.25 94.85 78.97 70.00 97.95 96.90 74.57 62.93 92.75 90.70
Average 56.24 52.34 74.97 75.75 69.31 59.29 94.73 94.08 72.58 63.91 97.23 96.71 70.72 60.05 91.52 90.50
Object Manipulation Carrier-v0 60.60 57.17 78.58 81.99 73.70 61.57 96.65 94.85 78.43 63.43 97.30 96.30 75.70 63.20 94.90 91.80
Carrier-v1 55.79 51.46 72.06 68.75 72.30 61.03 95.00 92.80 72.07 62.10 96.45 95.60 75.37 65.67 91.40 89.95
Pusher-v0 52.07 53.73 72.63 73.91 68.27 61.83 94.65 93.80 78.77 64.53 97.50 96.40 72.23 62.43 90.30 89.35
Pusher-v1 59.26 59.34 76.37 76.36 62.70 62.18 95.05 95.87 65.33 61.60 96.70 97.75 61.73 59.43 93.95 93.15
Average 56.93 55.43 74.91 75.25 69.24 61.65 95.34 94.33 73.65 62.92 96.99 96.51 71.23 62.68 92.64 91.06
Climbing & Balancing Climber-v0 54.59 53.33 85.29 85.67 58.83 58.60 96.15 95.90 59.47 58.10 97.70 97.00 58.57 58.63 92.35 93.85
Jumper-v0 9.70 33.42 91.36 81.32 47.97 50.70 93.25 91.99 39.83 49.00 96.20 95.95 27.27 41.90 93.20 90.90
Balancer-v0 63.38 60.96 78.72 79.46 69.10 68.26 96.45 95.66 68.27 71.47 97.40 98.20 71.97 69.17 91.85 92.80
Average 42.56 49.24 85.12 82.15 58.63 59.19 95.28 94.52 55.86 59.52 97.10 97.05 52.60 56.57 92.47 92.52
Average 55.34 52.88 79.28 77.58 66.62 60.05 95.07 94.27 68.75 62.48 97.12 96.73 66.37 60.06 92.13 91.19
Finding 1: Performance Hierarchy Among Models. The performance of the evaluated models exhibits a clear hierarchy across both accuracy and consistency metrics. Gemini-1.5-Pro demonstrates the highest accuracy in both easy tasks (68.75%) and hard tasks (62.48%), surpassing other models across various environments. This suggests that Gemini-1.5-Pro is more effective in identifying optimal robot designs across a range of task complexities. In contrast, Gemini-2.0-flash and Grok-2 exhibit moderate performance, with average easy accuracies of 66.62% and 66.37%, respectively, and hard accuracies around 60%, showing competitive but slightly lower capabilities compared to Gemini-1.5-Pro. Conversely, GPT-o3-mini exhibits the lowest accuracy (55.34% in easy tasks and 52.88% in hard tasks), with notable struggles in tasks like Jumper-v0 (9.70% easy), indicating challenges in generalizing effectively, particularly in complex scenarios requiring nuanced decision-making.
Finding 2: Trade-off Between Accuracy and Consistency. In this evaluation, there is no clear trade-off between accuracy and consistency; instead, performance in both metrics aligns with overall model capability. Gemini-1.5-Pro achieves the highest consistency (97.12% in easy tasks and 96.73% in hard tasks) alongside its top accuracy, highlighting its ability to maintain stable and reliable predictions across all task types. Gemini-2.0-flash follows with strong consistency (95.07% easy, 94.27% hard), outperforming Grok-2 (92.13% easy, 91.19% hard) despite similar accuracy levels, making it suitable for tasks where repeatability is critical. In contrast, GPT-o3-mini shows significantly lower consistency (79.28% easy, 77.58% hard) and accuracy, suggesting that its predictions are less stable and less reliable, particularly as task complexity increases.
Finding 3: Sensitivity to Task Difficulty. The performance of the models reveals sensitivity to task difficulty, with all models showing reduced precision in harder tasks where distinguishing between robot designs may be more challenging. For example, in Walker-v0, the precision drops from 75.40% (easy) to 65.20% (hard) for Gemini-1.5-Pro and from 57.65% (easy) to 56.58% (hard) for GPT-o3-mini, indicating that subtler performance differences in harder tasks pose difficulties. This trend is pronounced in tasks such as Jumper-v0, where the accuracy of GPT-o3-mini increases from 9. 70% (easy) to 33. 42% (hard), but remains low, suggesting an inconsistent sensitivity to task complexity. Gemini-1.5-Pro maintains relatively high performance across both easy and hard tasks, while Gemini-2.0-flash sustains strong consistency (e.g., 94.15% easy and 94.75% hard in Walker-v0), implying it relies on stable predictions even when accuracy decreases slightly. GPT-o3-mini struggles most, with low accuracy and consistency in complex tasks, highlighting its limitations in fine-grained evaluation.
Finding 4: Weaknesses of Models in Certain Environments The tasks Jumper-v0 and BidirectionalWalker-v0 reveal significant weaknesses in the evaluated model. All models exhibit very low accuracy, but high consistency, indicating a persistent inability to identify effective designs. This may be attributed to the specific characteristics and descriptions of the environments and tasks. For example, Jumper-v0 likely demands precise timing or control (e.g., jumping mechanics), which LLMs may not be adept at handling due to their lack of specialized reasoning for dynamic tasks. The bidirectional nature of the task BidirectionalWalker-v0 may also introduce ambiguous or conflicting cues (e.g., deciding between forward and backward movement), confusing the models about the correct course of action. The models’ architectures might not be suited to handle bidirectional decision-making, which requires understanding symmetric or reversible actions, leading to errors in design selection.
TABLE III: Global DWA metrics across different LLMs.
Environment GPT-o3-mini Gemini-2.0-Flash Gemini-1.5-Pro Grok-2
Walker-v0 58.68 65.52 67.82 64.71
BridgeWalker-v0 45.39 66.50 72.33 68.11
BidirectionalWalker-v0 52.11 55.55 54.31 54.58
UpStepper-v0 58.54 63.23 65.46 64.74
GapJumper-v0 57.11 64.38 74.10 68.94
Carrier-v0 60.07 66.92 69.80 69.08
Carrier-v1 53.84 66.02 66.55 70.65
Pusher-v0 53.01 63.81 68.77 65.80
Pusher-v1 60.53 62.19 63.69 60.77
Climber-v0 54.12 58.26 58.00 57.69
Jumper-v0 19.31 49.19 44.69 34.86
Balancer-v0 62.70 69.28 70.95 71.39
Average 52.96 61.99 63.72 61.66
Figure 6: Error distribution across difficulty levels for LLMs.
IV-C Performance Analysis Across Reward Difference Level
To assess the ability of LLMs to select better designs across varying levels of difficulty, we use the DWA metric, as introduced in Section IV-A. The results for these metrics are presented in Table III. Table III presents a comparative analysis of global Difficulty-Weighted Accuracy (DWA) metrics across four distinct LLMs: GPT-o3-mini, Gemini-2.0-Flash, Gemini-1.5-Pro, and Grok-2. The table breaks down the DWA performance for each LLM across twelve different EvoGym environments, providing a granular view of their design selection capabilities under varying task contexts. Notably, the DWA scores fluctuate across environments and LLMs, indicating varying degrees of success in discerning superior robot designs, weighted by the difficulty of each question. The ’Global Average’ row summarizes the overall performance of each model across all environments, offering a consolidated metric for comparison. Gemini-1.5-Pro emerges with the highest global average DWA of 63.72%, suggesting a marginally superior capability in embodied design reasoning within the RoboCrafter-QA benchmark compared to the other evaluated models.
Furthermore, we visualize the distribution of incorrect answers in different levels of reward difference in Figure 6. In particular, a significant proportion of errors made by LLM occur within the 0.8–1.0 difficulty interval, further highlighting their current limitations in making fine-grained distinctions between designs.
IV-D Ablation Study on Minor Prompt Design
We conduct an ablation study on prompt design to determine the key factors influencing the LLMs’ ability to make correct selections and assess how different information affects their performance. Given that the numerical representation of voxel materials is crucial for design, we focus our ablation study on task and actuator descriptions. Additionally, we perform an experiment where we modify the prompt to instruct the LLMs to select the Worse design instead of the better one, further analyzing the robustness of decision-making process of LLMs.
TABLE IV: Ablation study results on prompt design
GPT-o3-mini Gemini-2.0-flash Gemini-1.5-pro Grok-2
NoAct 54.09 66.93 69.78 65.82
NoEnv 52.08 62.47 62.50 62.91
Worse 51.93 58.48 67.87 61.24
Full 55.34 66.62 69.75 66.37
The ablation study on prompt design highlights the critical role of task and actuator descriptions in enabling language models (LLMs) to select optimal designs. The results in Table IV show that removing the task description (NoEnv) significantly reduces performance across all models, for example, from 55.34% to 52.08% for GPT-o3-mini and from 69.75% to 62.50% for Gemini-1.5-pro, underscoring its importance in guiding the decision-making process of LLMs. In contrast, omitting the actuator description (NoAct) has a minimal effect, with performance remaining stable or slightly varying. This may indicate that without actuator information, LLMs assume that the actuator is capable of maximizing the design’s reward. Additionally, when instructed to select the Worse design, the models exhibit lower accuracy compared to the full-information prompt (e.g., 66.62% to 58.45% for Gemini-2.0-flash), suggesting they are less adept at identifying inferior designs, likely due to a training bias toward selecting better ones. These findings emphasize the need for comprehensive task descriptions to maximize LLM performance in design selection tasks.
TABLE V: Comparison of design strategies.
Method Valid Design Reward Mean Reward Variance
Random 50/50 0.044 0.405
w/o Reference Knowledge 19/50 -0.607 0.255
w/ Reference Knowledge 50/50 0.115 0.035
IV-E Case Study of Robot Design initialization via LLM
In addition to selecting the optimal design in multiple-choice tasks, we aim to evaluate whether current large language models (LLMs) exhibit zero-shot capabilities for design initialization using reference information. Specifically, we use Walker-v0 and Jumper-v0 as reference environments, providing the LLM with 10 high-reward and 10 low-reward designs from each. The model is then tasked with generating a feasible structure for the GapJumper-v0 environment, leveraging environment details, $E=(D,T,C)$, along with the design space. We select Gemini-1.5-Pro as the best-performing model based on results from Table II and compare its performance against randomly generated designs and those produced without reference knowledge from other environments. The results are summarized in Table V.
With reference knowledge, all 50 designs were valid, yielding the highest mean reward of 0.115 and the lowest variance of 0.035, reflecting strong and consistent performance. Without reference knowledge, only 19 of 50 designs were valid, with a mean reward of -0.607 and a variance of 0.255. The random baseline achieved 50 valid designs but had a lower mean reward of 0.044 and a higher variance of 0.405, indicating poorer and less stable outcomes. These results show the model effectively transfers knowledge from reference environments to initialize designs in a new environment, demonstrating zero-shot design generation capabilities.
@@ V Conclusion
This paper introduces RoboCrafter-QA, a benchmark for evaluating multi-modal LLMs’ embodied design reasoning in soft robot design. Using the EvoGym simulator, we generated a diverse QA dataset capturing relationships between task requirements, environmental constraints, and robot morphologies. Experiments show that while LLMs perform well on simpler design selection, they struggle with subtle trade-offs and complex environments. RoboCrafter-QA provides a valuable tool for assessing and advancing LLMs’ design intelligence, enabling more automated and intelligent soft robot design.
Limitations and Future Work. This work has limitations suggesting future research directions. First, using pre-trained PPO policies could be improved by exploring LLM-driven control policy optimization. Second, the $5\times 5$ voxel grid restricts design complexity; a larger design space presents simulation challenges. Third, sim-to-real transfer, including material properties and controller transferability, is needed for real-world applications. Finally, multi-modal prompts (visual, tactile) could enhance LLMs’ embodied design understanding. Addressing these limitations will advance LLM-driven soft robot design.

### v2

@@ IV-B Benchmarking Results of Robot Design Selection
We evaluate the most advanced available LLMs across 12 diverse environments, using metrics categorized into ‘Easy’ and ‘Hard’ tasks in Table II, together with a heuristic baseline method.
Finding 1: Simple heuristics are insufficient. The actuator-count heuristic provides a non-trivial baseline (Avg. Acc: 47.07% Easy, 50.67% Hard), but its sub-51% accuracy confirms simple metrics are insufficient for this complex evaluation. Grok-4 (Avg. Acc: 59.72%/55.90%) and Gemini-2.5-Pro (52.12%/52.73%) consistently outperform this baseline, capturing more complex structure-function relationships. Conversely, GPT5’s accuracy (50.96%/50.46%) is on par with or worse than the heuristic, suggesting its high consistency stems from less effective internal biases.
Finding 2: A clear hierarchy of model capabilities and trade-offs. A distinct performance hierarchy emerges. Grok-4 shows the strongest reasoning, leading in average accuracy (59.72% Easy, 55.90% Hard). Conversely, GPT5 achieves the highest consistency (77.19%/77.42%) but its accuracy ($\approx$51%) is substantially lower, failing to reliably beat the heuristic. Gemini-2.5-Pro offers a balanced profile (Acc: $\approx$52%, Cons: $\approx$71%). Qwen-0.5-Instruct performs worst on both metrics, falling well below the baseline.
Finding 3: Task-dependent performance reveals systematic model biases. Model performance is highly task-dependent, revealing specific model biases. The Jumper-v0 task is exceptionally challenging: GPT5 and Gemini-2.5-Pro achieve minimal accuracy (9.84% and 16.84%, respectively). Critically, this failure is paired with high consistency (91.67% and 86.33%), indicating a convergence on a suboptimal design, not random guessing. This suggests a misinterpretation of dynamic propulsion. In contrast, strong performance on Carrier-v0 and UpStepper-v0 (65-69% accuracy) implies LLMs reason better about structured locomotion and balance than about impulsive dynamics.
IV-C Ablation Study on Minor Prompt Design
TABLE III: Ablation study results on prompt design.
Setting Qwen-0.5-Instruct GPT5 Gemini-2.5-Pro Grok-4
NoAct 53.31 47.62 50.79 54.72
NoEnv 51.31 46.96 49.79 56.72
Worse 26.61 42.66 47.45 58.38
Full 37.61 50.96 52.12 59.72
We conduct an ablation study on prompt design to examine how different components of the textual input affect each model’s ability to reason about robot design quality. Since voxel-based actuator encoding is essential for mechanical interpretation, we focus the ablations on task description and actuator specification, and additionally include a logical inversion setting where the models are instructed to select the worse design instead of the better one, to assess robustness in decision reversal.
Finding: Role of contextual cues and model capacity in prompt understanding. As shown in Table III, removing key contextual information leads to varied effects across models. For GPT-5, Gemini-2.5-Pro, and Grok-4, the full prompt yields the highest accuracy ($\approx$51–60%), confirming that complete task and environment cues are essential for consistent reasoning. In contrast, the smaller Qwen-0.5B shows unexpectedly higher performance when actuator or environment details are removed (NoAct: 53.31%), surpassing its own full-prompt accuracy (37.61%). This inversion likely stems from its limited capacity and overfitting to surface correlations, allowing it to perform better when the prompt is simplified. Across all models, the incomplete-prompt setting consistently degrades performance, indicating that logical negation and reasoning reversal remain challenging even for larger LLMs, which tend to retain a strong bias toward selecting optimal designs. Overall, these results highlight that task and environment context are indispensable for high-capacity models, while smaller LLMs may benefit from prompt simplification to avoid confusion or over-conditioning.
IV-D Finetuning LLM for Robot Design Selection
Figure 6: LLM finetuning pipeline for soft robot design.
TABLE IV: Finetuned Qwen-0.5B-Instruct performance across 12 environments for robot design selection.
Setting Accuracy Consistency
Easy Hard Easy Hard
Walker-v0 99.50 95.00 99.67 95.00
BridgeWalker-v0 100.00 96.00 100.00 97.00
BidirectionalWalker-v0 77.83 69.67 58.67 55.67
UpStepper-v0 97.00 85.50 95.33 79.00
GapJumper-v0 96.33 92.50 98.00 89.00
Carrier-v0 99.67 92.00 99.33 89.33
Carrier-v1 96.17 91.33 95.33 90.67
Pusher-v0 100.00 94.50 100.00 92.33
Pusher-v1 89.33 86.50 84.33 80.00
Climber-v0 86.83 83.17 80.00 77.33
Jumper-v0 98.17 89.50 97.33 87.67
Balancer-v0 87.33 84.00 79.67 81.00
Average 94.01 88.31 90.64 84.50
TABLE V: Average and best rewards across 10 generated robot designs in 3 real-world evaluated environments.
Setting GPT-5 Gemini2.5-Pro Grok-4 Ours
Avg Best Avg Best Avg Best Avg Best
Walker-v0 0.35 0.86 0.17 0.43 0.02 0.64 2.98 4.51
Pusher-v0 0.15 0.57 0.20 0.73 0.01 0.10 0.69 2.64
Carrier-v0 -0.36 -0.22 -0.26 -0.04 -0.53 -0.37 0.79 1.90
Average 0.05 0.40 0.04 0.37 -0.17 0.12 1.48 3.02
To unleash the full potential of LLM model’s understanding the structure of soft modular robot, we finetune Qwen2.5-0.5B-Instruct using a parameter-efficient Low-Rank Adaptation (LoRA) strategy [34]. As illustrated in Fig. 6, all training data are organized by environment and difficulty, where each folder contains JSON files of question–answer (QA) pairs. Each sample is formatted in the Qwen chat style, consisting of a system prompt (‘You are a helpful assistant that answers multiple choice questions accurately’), a user query describing the environment, task, and structure design, and an assistant response in the form ‘The answer is <answer>.’ The complete sequence, which combining system, user, and assistant messages, is tokenized and right-padded to a fixed length before being passed into the transformer.
During finetuning, LoRA adapters are injected into both the attention and feed-forward projections, specifically the layers q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj, while all other parameters remain frozen. The model is optimized under a causal language modeling objective that predicts the assistant’s next token given all preceding context. Gradients are propagated only through the LoRA components, allowing the model to efficiently learn instruction-level reasoning for accurate evaluation of soft-robot designs across diverse environments.
Figure 7: A real-world replica of soft modular robot designs in simulation. (a) Real modular soft robot with (b) four basic modules assembled via tiny magnets integrated in each module. Real-world speed comparison of modular soft robot designs for three tasks: (c) Walker-v0, (d) Pusher-v0, and (e) Carrier-v0. Each task compares two modular designs from Table VI.
The results of this finetuning process, detailed in Table IV, demonstrate that this approach achieves exceptionally strong performance. The finetuned 0.5B model reaches a high average accuracy of 94.01% on easy tasks and 88.31% on hard tasks across all 12 environments. This high accuracy is complemented by strong average consistency of 90.64% (Easy) and 84.50% (Hard). Notably, the model achieves perfect 100% accuracy and consistency on easy tasks for environments like BridgeWalker-v0 and Pusher-v0. This indicates that our parameter-efficient finetuning strategy is highly effective at enabling even a small model to understand and accurately evaluate complex robotic designs.
TABLE VI: Comparison of simulation rewards and real-world metrics for robot designs across three tasks.
[…]
Walker-v0 Design A 3.17 1.93
Design B 3.26 3.25
Pusher-v0 Design A 2.25 3.54
Design B 2.78 4.35
Carrier-v0 Design A 0.95 2.26
Design B 1.32 2.92
IV-E LLM for Direct Robot Design
After finetuning, we further evaluated the model’s generative capability by prompting it to directly design soft-robot structures instead of merely selecting between options. To ensure manufacturability and real-world verification in Sec. IV-F, we constrained the LLM to generate designs containing no more than five actuators across three selective environments. As shown in Table V, the finetuned Qwen-0.5B model (“Ours”) achieved the highest average and best rewards across all three environments: Walker-v0, Pusher-v0, and Carrier-v0, which largely outperforms existing general-purpose LLMs, including GPT-5, Gemini-2.5-Pro, and Grok-4. Our model obtained a mean average reward of 1.48, exceeding the next best baseline by over 10×, and a mean best reward of 3.02, demonstrating that the finetuned model not only understands design-quality distinction but can also generate effective and high-reward robot morphologies.
These results confirm that instruction-level finetuning with structured QA data effectively transfers to real design generation, enabling compact LLMs to reason about and synthesize performant soft-robot structures across environments.
IV-F Real-world Robot Design Verification
To validate our simulated findings, we established a pipeline for fabricating and testing physical voxel-based soft robots that mimic our simulation setup (Fig. 7a-b). We directly mirrored the $\times5$ grid representation used in our simulations, ensuring a one-to-one mapping between the digital design and its physical counterpart. Each cell in the $\times5$ design matrix was physically instantiated as its corresponding component: a soft voxel, a rigid voxel, a horizontal actuator, a vertical actuator, or a void. All components were fabricated using casting molds. We selected materials based on function: highly elastic Ecoflex™ 00-50 (Smooth-On, Inc; Shore 00-50) was used for the soft voxels and actuator bodies, while Dragon Skin™ 30 (Smooth-On, Inc; Shore 30A) was used for the rigid voxels. To maintain linear expansion while constraining motion in other directions, the pneumatic actuators were fabricated with embedded strings wound around the elastomer.
For modular assembly, magnets were installed at the vertices of the soft modules, enabling rapid and reconfigurable connections between voxels. The assembled physical robots are driven by a pneumatic actuation system. The inflation and deflation of each actuator are governed by a three-way valve, which is managed by a PWM-based controller. This complete actuation and control setup allows for temporal control over the robot’s gait and movement, enabling direct comparison with our simulated results. To better align with the simulation hyperparameters, particularly regarding the coefficient of friction that influences motion, we performed some engineering optimizations in the real-world experiments. Specifically, we added foot-like structures at the robot–ground contact points with anisotropic friction and introduced a testing surface with specific frictional properties. These modifications allowed the robot to utilize ground friction more effectively for forward motion.
Our real-world experiments, summarized in Table VI and illustrated in Fig. 7c-e, successfully confirmed a strong alignment between simulated performance and physical reality. The designs used in the experiments were generated by our finetuned LLM, including those with higher simulated rewards and those with lower rewards. The results demonstrate a clear positive correlation: designs that achieved higher rewards in simulation consistently translated to higher real-world performance, as measured by a faster forward speed.
@@ V Conclusion
This work demonstrates that while generalist LLMs fail at nuanced, embodied design reasoning, this capability can be unlocked through targeted training. We introduced RoboCrafter-QA, a benchmark for soft robot design, which first revealed the limitations of SOTA models. We then used this benchmark to finetune an efficient LLM that achieves SOTA design selection and successfully generates novel, high-performing robot morphologies. By fabricating and testing these designs on a physical modular robot, we validated a strong sim-to-real correlation. This work provides a complete pipeline, proving that LLMs can be specialized into effective co-designers for real-world physical systems.
Limitations and Future Work. First, using pre-trained PPO policies could be improved by exploring LLM-driven control policy optimization. Second, the $5\times 5$ voxel grid restricts design complexity; a larger design space presents simulation challenges. Third, sim-to-real transfer, including material properties and controller transferability, is needed for more robust real-world applications. Finally, multi-modal prompts (e.g., visual prompts) could further enhance the robot design capability of foundation models.